Thomas’ Calculus Early Transcendentals 13ed

Homework Assignment for Chapter 07

1. Section 7.2: problems 9, 13, 17, 21.

2. Section 7.3: Problems 1, 11, 15, 21, 37, 45, 47, 53, 57.

3. Section 7.4: Problems 3, 5, 7, 8, 9, 10, 23(a), 24.